OpenStudy (anonymous):

@iGreen @Indogot find the sum 5 9/13+ 5 9/13 A.13/74 B.10 12/13 C.1 15/13 D. 13/142

OpenStudy (anonymous):

I got 19 5/13 but i dont see that up there...

OpenStudy (anonymous):

nvm i saw the equation wrong give me a sec

OpenStudy (igreen):

Add the whole numbers and add the fractions. \(5 + 5 =~?\) \(\dfrac{9}{13} + \dfrac{9}{13} \rightarrow \dfrac{9+9}{13}\)

OpenStudy (anonymous):

18

OpenStudy (anonymous):

ok heres what i got: 11 5/13

OpenStudy (anonymous):

IS THAT RIGHT

OpenStudy (igreen):

@J-bird Don't give answers.

OpenStudy (igreen):

Yes, 9 + 9 is 18..so that gives you 18/13. What's 5 + 5?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Which gives you 10 18/13.

OpenStudy (igreen):

Yep, that gives you 10 18/13, but 18/13 is an improper fraction so you have to change it. What's 18 - 13? @msdf!!!!!